# DeepAndes: A Self-Supervised Vision Foundation Model for Multi-Spectral Remote Sensing Imagery of the Andes

### v1

@@ VI Discussion
The evaluation of downstream tasks indicates that DeepAndes enhances visual feature representations, which improve as pre-training scales increase. With the 3-million scale pre-training, DeepAndes achieves over 83% accuracy in few-shot imbalance classification, surpassing other models, including those trained from scratch in a supervised manner, across various fine-tuning scales. Similar patterns are evident in image retrieval and segmentation tasks, where the feature embeddings from the pre-trained backbone show potential for both image-level and pixel-level remote sensing tasks when labeled data is limited.
@@
There are also some limitations to this work. First, the data for self-supervised pre-training should undergo more thorough data curation. In [36], the training data undergo a comprehensive data de-duplication process and employ image retrieval to refine the uncurated data source with curated sources such as ImageNet-22 [14]. In our work, we try to sample from diverse archaeological regions while scaling up the datasets, yet we lack curated datasets and detailed prior knowledge of the survey regions, unlike established benchmarks. The archaeological regions are not as distinct as these natural imaging datasets, suggesting a need for further de-duplication beyond excluding featureless samples. Additionally, as we refine the DeepAndes model, it could be utilized to analyze and help reduce duplicate samples in the pre-training dataset.
@@
For the future and in progress work, the DeepAndes will be integrated into the the GeoPACHA geospatial platform [46] as a web application, collaborating regional experts and their teams. Meanwhile, depending on the remote sensing tasks, the human-in-the-loop is also included as experts will also verify and compare the autonomous model prediction with their data annotations. The curated data will be used for next round of foundation model fine-tuning to acquire more a specialized model, DeepAndesArch, for remote sensing tasks. From a model perspective, the proposed foundation model offers broad utility and can be further fine-tuned for advanced object detection tasks. By incorporating complex transformer heads, such as Co-DETR [54], it can provide valuable support to archaeological survey teams.
@@ VII Conclusion
In this work, we present DeepAndes, the first AI vision foundation model for multi-spectral remote sensing data for social and earth science applications, utilizing the SOTA DINOv2 framework. DeepAndes is pre-trained on 3 million WorldView-2 and WorldView-3 satellite images with eight spectral bands. Through extensive downstream experiments on three prevalent computer vision tasks in archaeology—imbalanced image classification, image instance retrieval, and semantic segmentation—DeepAndes demonstrates its effectiveness in both image-level and pixel-level feature representation, as well as in few-shot learning capabilities. The pre-trained DeepAndes will be integrated into the GeoPACHA web app to expand the scale of our archaeological surveys in the Andes with human-in-the-loop verification. The broad utility of the proposed foundation model can be further fine-tuned for more advanced object detection tasks using complex transformer heads, such as Co-DETR. As we collect more data with the AI-assisted GeoPACHA tool, experts will be able to contribute more effectively as observers and analysts in the archaeological workflow.
@@ Acknowledgements
This work was supported by the following grants: National Endowment for the Humanities Level III Digital Enhancement Grant (Award ID Number HAA-293452-23), National Science Foundation IIS-III: Medium: Collaborative Research (Award number 2106717); National Science Foundation Collaborative Research: Research Infrastructure: HNDS-I (Award Numbers 2419793 and 2419794); Vanderbilt University Scaling Success Grant, Vanderbilt University Discovery Grant. The project also benefited from the infrastructural support of the Vanderbilt University Data Science Institute (GPU computational infrastructure) and the Vanderbilt University HLRB Lab (computational infrastructure), and the Spatial Analysis Research Laboratory (geospatial computational infrastructure).

### v2

@@ VI Discussion
The evaluation of downstream tasks indicates that DeepAndes enhances visual feature representations, which improve as pre-training scales increase. With the 3-million scale pre-training, DeepAndes achieves over 83% accuracy in few-shot imbalance classification, surpassing other models, including those trained from scratch in a supervised manner, across various fine-tuning scales. Similar patterns are evident in image retrieval and segmentation tasks, where the feature embeddings from the pre-trained backbone show potential for both image-level and pixel-level remote sensing tasks when labeled data is limited. In addition to the main findings, it is important to consider the following aspects.
VI-A Limitations
@@
Limited Pre-training and Data Curation. A limitation of this work is that the data for self-supervised pre-training should undergo more thorough curation. In [38], the training data undergo a comprehensive data de-duplication process and employ image retrieval to refine the uncurated data source with curated sources such as ImageNet-22 [14]. In our work, we try to sample from diverse archaeological regions while scaling up the datasets, yet we lack curated datasets and detailed prior knowledge of the survey regions, unlike established benchmarks. The archaeological regions are not as distinct as these natural imaging datasets, suggesting a need for further de-duplication beyond excluding featureless samples.
Inferior Experimental Outcomes. Figure 2 (scaling law) shows that larger-scale pre-training yields greater benefits when labeled data are scarce. However, limitations remain. In (1) imbalanced classification (Table II) with only 72 positive samples, performance gaps persist across pre-training scales (30K, 300K, 3M). In some contexts, obtaining millions of samples for pre-training can be difficult, and downstream tasks may contain even fewer positive samples, highlighting the need for improved pre-training and fine-tuning with imbalance-aware methods. For (2) few-shot segmentation, results show considerable noise (e.g., Figure 7, middle panel). Small foreground objects are particularly difficult, suggesting the object detection task may be more effective. Distinguishing active from archaeological corrals is also challenging when weathering obscures animal evidence, suggesting the integration of expert domain knowledge (e.g., elevation, coordinates, or textual data) to domain fine-tuning. For (3) image-level content retrieval, input queries with many loci types can yield noisy results, as features of archaeological interest may be overshadowed by others; pixel- or object-level retrieval will be needed in future work.
Feature Scaling Challenge. Similar to other Earth observation (EO) studies, features in our data vary greatly in scale, from 10-meter archaeological structures to corrals spanning hundreds of meters. As shown in the qualitative results of few-shot loci segmentation (Figure 7), using a 224-pixel input (the default for natural image benchmarks and many existing EO models) poses challenges: buildings appear as very small objects, while corrals may exceed the image crop size. In real archaeological surveys, researchers often switch between local and large-scale views to help them identify features of varying sizes and distributions. This suggests incorporating larger image crop sizes—i.e., higher pixel dimensions—during pre-training and specific fine-tuning. Larger image crop size can capture richer semantic information, potentially benefiting a wider range of downstream applications.
VI-B Broad Impacts
Value to Archaeology. Although costly, developing a foundation model can greatly improve our ability to identify archaeological features in high resolution satellite imagery, given the variability of landforms and archaeological features across the Andes. The few-shot results in classification and segmentation demonstrate a scaling law with increasing pre-training scale, confirming the effectiveness of the SSL framework. This framework provides a foundation model for diverse downstream tasks while reducing the burden of retraining and large-scale labeling. Additionally, through zero-shot image retrieval, the pre-trained backbone can rapidly mine and expand relevant datasets for specific archaeological objectives, a task infeasible at scale with manual methods.
Domain Generalization. While the proposed foundation model is explicitly pre-trained on extensive high-resolution satellite imagery from the Andes, its potential applications extend beyond this region. For archaeological patterns unique to the Andes, iterative fine-tuning with remote sensing imagery from additional areas could further enhance the model’s generalizability across a broader global span, supporting a wider range of computational applications. Beyond archaeology, the model may also advance geoscientific inquiries, including studies of land use, human–environment interaction, and cultural heritage preservation. In this sense, the Andes serve as a proving ground, demonstrating how large-scale pre-training can be well-suited for domain-specific applications such as archaeology.
VI-C Future Work
@@
For the future and in progress work, the DeepAndes will be integrated into the the GeoPACHA geospatial platform [49] as a web application, collaborating regional experts and their teams. Meanwhile, depending on the remote sensing tasks, the human-in-the-loop is also included as experts will also verify and compare the autonomous model prediction with their data annotations. The curated data will be used for next round of foundation model fine-tuning to acquire more a specialized model, DeepAndesArch, for remote sensing tasks. From a model perspective, the proposed foundation model offers broad utility and can be further fine-tuned for supporting more advanced archaeological remote sensing tasks.
@@ VII Conclusion
DeepAndes is the first AI vision foundation model for multi-spectral remote sensing data for social and earth science applications, utilizing the SOTA DINOv2 framework. DeepAndes is pre-trained on 3 million WorldView-2 and WorldView-3 satellite images with eight spectral bands. Through extensive downstream experiments on three prevalent computer vision tasks in archaeology—imbalanced image classification, image instance retrieval, and semantic segmentation. We demonstrates the effectiveness of DeepAndes for both image-level and pixel-level feature representation, as well as in few-shot learning capabilities. The pre-trained backbone will be integrated into the GeoPACHA web app to expand the scale of our archaeological surveys in the Andes with human-in-the-loop verification. The broad utility of the proposed foundation model can be further fine-tuned for more advanced archaeological remote sensing tasks. As we collect more data with the AI-assisted GeoPACHA tool, experts will be able to contribute more effectively as observers and analysts in the archaeological workflow.
@@ Acknowledgements
This work was supported by the following grants: National Endowment for the Humanities Level III Digital Enhancement Grant (Award ID Number HAA-293452-23), National Science Foundation IIS-III: Medium: Collaborative Research (Award number 2106717); National Science Foundation Collaborative Research: Research Infrastructure: HNDS-I (Award Numbers 2419793 and 2419794); Vanderbilt University Scaling Success Grant, Vanderbilt University Discovery Grant. The project also benefited from the infrastructural support of the Vanderbilt University Data Science Institute (GPU computational infrastructure) and the Vanderbilt University HLRB Lab (computational infrastructure), and the Spatial Analysis Research Laboratory (geospatial computational infrastructure). This manuscript has been co-authored by UT-Battelle LLC, under contract DE-AC05-00OR22725, with the Oak Ridge National Laboratory (ORNL) of the US Department of Energy (DOE). The U.S. Government retains a nonexclusive,worldwide license to publish or reproduce the published form of this manuscript, or to authorize others to do so, for U.S.Government purposes, as acknowledged by the publisher.